OpenStudy (anonymous):

Implicitly differentiate 3y + 1 = (y - 1) x + y^2 (x - 2)

OpenStudy (anonymous):

\[3y + 1 = (y - 1) x + y^2 (x - 2)\]

OpenStudy (anonymous):

Distribute the RHS to get:\[3y + 1 = yx - x + y^2x - 2y^2\]Differentiate both sides with respect to x. Remember the product rule, we are left with:\[3y \prime + 1 = y - y \prime x - 1 + y^2 + 2yy \prime x - 4yy \prime\]

OpenStudy (anonymous):

Oops, without the +1 on the LHS.

OpenStudy (anonymous):

Separate all terms with y' on the LHS and we get:\[y \prime (3 + 4y + x - 2yx) = y - 1 + y^2\]Divide both sides by (3 + 4y + x - 2yx) and we arrive at:\[y \prime = \frac{y^2 + y - 1}{-2yx + x + 4y + 3}\]

OpenStudy (anonymous):

Small sign error, it should be:\[ y \prime = - \frac{y^2+y−1}{2xy+x−4y−3}\]

OpenStudy (anonymous):

ok got close justa few sign errors myself... but is that all I feel like I am missing something at the end

OpenStudy (anonymous):

Nope, that's the answer, unless you want to factor out y on top or try to do some algebra. The y' is a function of (y,x) when defined implicitly.

OpenStudy (anonymous):

ok well thanks you were a huge help!!!

OpenStudy (anonymous):

No problem :-)